Please help me solve this inequality
!4x-3!<5
The < has a line below it...

Answer by edjones(8007)

!4x-3!<=5
.
4x-3<=5
4x<=8
x<=2
.
4x-3>=-5
4x>=-2
x>=-1/2
.
Ed